Journalist Report – February 18th

Crew 205 Crew Journalist Report 18-FEB-2019 Sol 8

Author: Dave Masaitis

Title: A Sol of Great Mystery

Sol 8 began as others have, with crew members emerging from quarters to start a tea kettle and make breakfast. Crew 205 quickly became inquisitive as they peered outside the portholes of our home to see a mysterious white precipitate covering the regolith ridges, with more falling from the sky. While it seemed to resemble snow on Earth, our previous understanding of Mars told us that the planet does not have enough water for such a weather phenomenon to be plausible. It was therefore imperative to treat this precipitate with suspicious caution, especially as it seemed to saturate the regolith and turn it to a rather slippery mud-type substrate. In the interests of crew safety, Commander Grulich reported to Mission Support and announced that our planned EVA would have to wait for another Sol.

Over breakfast, Commander Grulich and Executive Officer Alotaibi consulted with the crew to realign work priorities for a Sol at the Hab. The mysteries of Sol 8 did not stop there. GreenHab Officer Nathan Hadland was analyzing rock and substrate samples in the science dome, when he called on the radio for the crew to come look at something. He had found some samples that, when washed, appeared to be the fossilized shells of some form of invertebrate. Had the crew stumbled on an ancient sea bed while on EVA? By visual appearance, these certainly appeared to be fossilized shells but further testing would be required before any confirmations could be made. Given that some invertebrates on Earth make their shells from calcite, it seemed reasonable that testing a fossilized fragment with a diluted hydrochloric acid solution should create a reaction…and it did! Samples taken from a sampling site within a couple hundred meters of the fossilized shells rendered beautiful crystalline structures, which are suspected to be either halite or gypsum. Both of these pieces of evidence point initially to a region of planetary surface that may have once been an ancient sea bed, so further investigation will be required. While it is far too premature of a discovery to draw any firm conclusions, the mysteries of todays analysis prove even more interesting than the mysterious white precipitate falling from the sky. Are the two perhaps connected?

Some analysis of the fossils may continue into the evening, and the crew will enter its communications window to file their daily reports to Mission Support. Following this, the Commander and Executive Officer will discuss the crew’s priorities of work for Sol 9, before having a dinner of salmon and hushpuppies prepared by one of our more apt cooks, Communications Officer Hannah Blackburn. After dinner, we will likely engage in some table games or perhaps watch Apollo 13, before laying down to get some rest before another Sol.

Operations Report – February 18th

Crew 205 Operations Report 18 – FEB – 2019

SOL: 8

Name of person filing report: Verónica Triviño

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on at 1535 (18thFeb), turned off at 1100 (18thFeb)

Solar— SOC % – Turned off Gen (18th Feb) 95%, Turned on Gen (18th Feb) 76%

Diesel Reading – 80%

Propane Reading – ~ 40%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 50%, ~ 276 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 205.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01413342 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 51.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 75.0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 75.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: N/A

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Yes, to town.

Crew Car used and why, where? N/A

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Screws tightened in the back airlock door. White tunnel covers were re-secured with new zipties, as well as tightening loose zipties and clipping the ends of the zipties.

Summary of Green Hab operations: Loft tank water filter replaced.

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: We still need a 75W standard light bulb.

Greenhab Report – February 18th

Crew 205 Green Hab Report – 18-FEB-2019
GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (0 hrs)
Shade Cloth (40%)
Average temperatures:

Low: 17.7°C

High: 23.5°C
Hours of supplemental light: 5 Hours

Daily water usage for crops: 11 gallons

Water in Blue Tank: ~ 205.5 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 09.40 AM 21.4° C / 31 %
2) 12.30 PM 19.2° C / 37 %
3) 15:30 PM 22.6° C / 47 %
4) 18:00 PM 20.6° C / 42 %

Changes to crops: Basil plants flowering and the “Mystery Greens” from the seed bank began sprouting. One of the remaining pots of cucumbers was scrapped.

Narrative: The Arabidopsis thaliana experiments are performing very well compared to the previous trial. Two regolith samples appear to be standing out as more viable candidates for plant growth substrates. The pH study on the regolith samples was completed today and the samples’ inability to support Arabidopsismay be in part due to their wettability properties and alkaline nature.

Miracle Grow nutrient supplement was sprayed sparingly on tomatoes, sugar snap peas, and basil. One of the remaining struggling cucumber plants died and the dead biomass weight will be used in our cellulosic ethanol ISRU project. The door remained closed throughout the day due to the consistent snowfall and consequently caused a fairly constant temperature throughout the day. 9.8 g of parsley and 6.4 g of basil were harvested as a garnish for lunch and dinner. Two of the basil plants flowered, which we groomed as a preventative measure against the plant entering another stage in its lifecycle. Upon further investigation, it appears that the radish plant that flowered is healthy contrary to our hypothesis in our last report.

Harvest: 9.8 g parsley

6.4 g of basil

Support/supplies needed: Several pots and the large grow bed appear to be over-saturated. Should we decrease the frequency of watering from 4 times a day to allow the soil to aerate.

Astronomy Report – February 18th

Crew 205 Astronomy Report 18 Feb 2019

Crew Astronomer:Natalia Larrea

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-WF

Objects Viewed:

No observations for MDRS Target 1 were captured using MDRS-14 last night due to cloudy weather. We are still waiting for the sky to clear to obtain further measurements.

This morning we submitted an additional observation request for Orion Nebula using MDRS-WF telescope. Observations were set at 3 sec exposure using RGB and Lum filters. Right Ascension correction for this object was set at -10 arcmin.

Problems Encountered: Bad weather conditions.

Sol Summary 8 – February 18th

CREW 205 Sol Summary

Sol: 8

Sol Summary title: Fossils, mystery and other things

Author: Maria Grulich

Mission status: Nominal

Sol activity summary: The EVA suits have been fixed last night successfully.

This morning the crew woke up with the smell of pancakes with blueberry and apple toppings prepared by commander Maria with a banana milkshake to top it off.

The planned EVA had a chance of 30% snow but when we looked outside the window, we just could see white everywhere. Therefore, the decision was made to cancel the EVA.

It gave us the chance to catch up with science and maintenance of the habitat. The crew astronomer for the day is Natalia and she submitted a new request with the help of Ghanim who is today’s XO, for the Orion nebula as yesterday no new observation could be made.

Veronica and Dave exchanged the water filter. The tunnels were brushed to clean them from snow. Nathan continued the characterization of the samples. He made some interesting observation as some of the samples that were taken yesterday turned out to be fossilized shells which is very exciting as this means it is a sign of a former sea at the Moon area. In the afternoon he took care of the ISRU plants in the Greenhab which luckily survived the night.

Dave prepared some noodle soup and tonight we are getting salmon and hushpuppies and rice by comms officer Hannah.

Besides that, the crew edited some of the videos and took some extra images for the outreach project.

The crew journalist of today took some nice pictures of the hab in snow and the fossil shells.

After dinner, we’re hopefully going to play some games or watch a movie as a team.

Look ahead plan: The crew intends to conduct the EVA planned for today again, for tomorrow if the weather outside allows it. In addition, the science continues with characterization and observation of the Arabidopsisseedlings and their development in the regolith substrate and the editing of the videos will continue.

Anomalies: EVA cancelled

Weather: snow all day

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Journalist Report, Astronomy Report

Operations Report – February 17th

SOL: 7

Name of person filing report: Verónica Triviño

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on at 1600 (17th Feb), turned off at 0830 (17th Feb)

Solar— SOC % – Turned off Gen (17th Feb) 82%, Turned on Gen (17th Feb) 100%

Diesel Reading – 85%

Propane Reading – ~ 40%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 55%, ~ 301 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 216.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01413047 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 51.0 hours

Beginning charge: 100%

Ending charge: 28%

Currently charging: Yes

Curiosity rover used: Used

Hours: 75.0 hours

Beginning charge: 100%

Ending charge: 44%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 75.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: N/A

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? N/A

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Some suits are being repaired in the evening. The parts of the ones repaired yesterday evening worked well on this morning’s EVA.

Summary of Hab operations: Screws tightened in the back airlock door. White tunnel covers were re-secured with new zip ties, as well as tightening loose zipties and clipping the ends of the zip ties.

Summary of Green Hab operations: GreenHab door screws tightened.

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report.

Journalist Report – February 17th

Author: Natalia Larrea

Title: “Oh I will survive!”…. with the taste of cinnamon rolls and the sound of “Starman”

We woke up with the smell of fresh coffee and home-made cinnamon rolls made by our COMMS Officer (Hannah). Thanks to her and her excellent culinary skills, we are surviving in the Red Planet eating food as if we were back home on Earth (and even better I would say!). All the Crew sat down to enjoy breakfast all together as usual, but this time listening to the sound of Gloria Gaynor and her “I will survive” and David Bowie’s “Starman” from Dave’s (HSO) iPad. Full of energy, all the crew and our new Martian mascot “Gus” (found a couple of days ago on the Martian surface) started singing and dancing to the rhythm of these songs with big smiles on our faces. It was a great way to start our working day.

In the morning, the EVA team composed by Daniel (Commander), Veronica (ENG), Nathan (GHO), and Dave (HSO) left at 10.00 am driving Curiosity and Opportunity heading to the “Moon Area”. The team continued collecting geological samples over there, following the initial research performed a few days ago. They made further discoveries including some fossil shells and interesting igneous rocks with crystals hidden inside. Meanwhile, in the Hab, Maria (XO) supervised overall workflow and conducted sample inventory in the GreenHab and in the Science Dome. In the afternoon, we could see some clouds rising in the horizon. By 1.30 pm some snowflakes started hitting the ground. Since “Crew safety and security first” is our main motto in this mission, the team decided to cancel the EVA scheduled for the afternoon. We took the opportunity to continue working on our different projects and tidy the Hab. Nathan (GHO) spent the afternoon in the GreenHab and the Science Dome doing some regolith sample characterization, watering the plants, and taking care of the A. Thaliana -ISRUexperiment (it is truly exciting to see how some of the seeds are growing in Mars like soil!). Hannah (COMMS) worked on food inventory and other tasks while Satana’s music was playing on the background. Dave (HSO) and Natalia (MEDIA) had the chance learn more about astronomy and variable stars thanks to the explanations from Ghanim (ASTRO). Martian mascot “Gus” is also helping us around the base. He is assisting us in putting together some outreach videos – dedicated to all of you Earthlings out there following us – about our life in here and about Mars (to be released soon… stay tuned!). This afternoon was also a time for doing some fixing around the base. Vero (ENG), Dave (HSO) tighten some loose screws here and there and together with Maria (XO) they fixed some loose sheets in the tunnels tighten all the cable ties and discovered that one bigger sheet is lose over the science dome tunnel. Our leadership team for the day – Daniel (Commander) and Maria (XO) – finished the evening working on the daily reports and planning for tomorrow.

Today we completed Sol number 7. This has been an important milestone as it marks the middle of our mission. It has been one week since we started exploring these lands, adapting to this new life, overcoming daily challenges, and advancing admirably in our research projects. EVA suits now feel like a further extension of us. We have become much more efficient in our daily operations thanks to the new routine and processes that we have implemented together as a team. All the daily tasks around the base seem now to be part of our normal life. But above all, it has been one week since we became a “Martian family”. We can feel now how we have become a single unified team, driven by our shared-passion of conducting science and exploration, and supporting each other in all the steps along the way to achieve our mission goals. We have found a great balance between hard work and relax time, pushing our personal boundaries and advancing in our mission objectives. We have passed already one week here and we cannot wait to see what excitements the next days will bring!

Ad Astra!

Astronomy Report – February 17th

Crew Astronomer: Ghanim Alotaibi

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-14

Objects Viewed: Last night 10 images were captured out of the requested 18 images for MDRS Target 1. The quality of the images is bad and one of the images looks shaky. Crew astronomer suggested to resubmit the observation and Observatory Director recommended the same due to some problems encountered, thus the observation was resubmitted.

New photometry measurement was performed for SY MON target. Based on the previous experience gained, the new measurement was found to be 14.132 with an error of 0.0185. The data point was submitted to the AAVSO website with the star 122 as a check star.

Problems Encountered: Mount Problems which caused shaky images. The moon was also bright, so bad images were captured. Observatory Director solved the problem.

Operations Report – February 17th

SOL: 7

Name of person filing report: Verónica Triviño

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on at 1600 (17th Feb), turned off at 0830 (17th Feb)

Solar— SOC % – Turned off Gen (17th Feb) 82%, Turned on Gen (17th Feb) 100%

Diesel Reading – 90%

Propane Reading – ~ 40%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 62%, ~ 343 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 216.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01413047 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 51.0 hours

Beginning charge: 100%

Ending charge: 28%

Currently charging: Yes

Curiosity rover used: Used

Hours: 75.0 hours

Beginning charge: 100%

Ending charge: 44%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 75.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: N/A

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? N/A

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Some suits are being repaired in the evening. The parts of the ones repaired yesterday evening worked well on this morning’s EVA.

Summary of Hab operations: Screws tightened in the back airlock door. White tunnel covers were re-secured with new zip ties, as well as tightening loose zipties and clipping the ends of the zip ties.

Summary of Green Hab operations: GreenHab door screws tightened.

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report.

GreenHab Report – February 17th

GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (~0.5 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 18.6°C

High: 26.3°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 216.5 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 09.25 AM 22.3° C / 23 %
2) 12.05 PM 24.4° C / 24 %
3) 15:15 PM 22.6° C / 69 %
4) 18:00 PM 21.5° C / 45 %

Changes to crops: Radish plant is flowering.

Narrative: The Arabidopsis thaliana experiments died overnight despite ample watering due to dehydration. Consequently, the experiments were restarted this afternoon (see Science Report). Miracle Grow nutrient supplement was sprayed sparingly on sugar snap peas, tomatoes, mustard, and cucumbers. Dead biomass on the sugar snap peas was pruned in order to help improve the productivity and health of the plants. The dead biomass weight will be used in our cellulosic ethanol ISRU project. The door was opened at 1515 hours to allow for ambient airflow but closed at 1545 because of the snow storm and drop in temperature. 4.2 g of Mixed Salad Greens were harvested as a garnish for dinner. One of the radish plants is flowering which may be cause for concern (angiosperms often flower when stressed). We will investigate when the radishes were planted in the logbook to determine whether this is part of the natural lifecycle or actually a stressed organism.

The GreenHab inventory and inventory of plants currently growing at the facility was completed today by the GreenHab officer and today’s XO. These will be made available for future crews and updated by crews during their rotation.

Harvest: Mixed salad greens: 4.2 g

Support/supplies needed: None.